A CNC machine centre costs £100,000. Its value depreciates at X% per annum.

(a) What will be its value after Y years?

(b) How long will it take to have a value of £50,000?



1
Expert's answer
2021-10-21T12:07:25-0400

a)

value after 1 year:

"100000-100000x=100000(1-x)"

after 2 years:

"100000(1-x)-100000(1-x)x=100000(1-x)(1-x)"

after y years:

"100000(1-x)^y"


b)

"100000(1-x)^y=50000"

"y=\frac{1}{2ln(1-x)}"